a student runs 2 laps around a track in 120 seconds. The distance around the track is 400 meters. What is the student's velocity for the run? what is the student's average speed forthe trip?
1
Expert's answer
2012-11-16T05:14:07-0500
a student runs 2 laps around a track in 120 seconds. The distance around the track is 400 meters. What is the student's velocity for the run? what is the student's average speed forthe trip?

The distance covered by running student is

D = 2*400[m] = 800[m].

So, the student's average speed for the trip is

S = D/T = 800[m]/120[s] = 20/3[m/s] = 6.(6)[m/s].